OpenStudy (anonymous):

If y varies inversely as x, and y = 12 as x = 2, find y for the x-value of 6 a.6 b.5 c.4 d.3

OpenStudy (dtan5457):

Formula for inverse variation is y=k/x y=12,x=2 solving for k.. 12=k/2 k=24 y=24/6 y=?

OpenStudy (anonymous):

Okay thank you!